### v1

# Risk-Budgeted Control Framework for Balanced Performance and Safety in Autonomous Vehicles

Abstract
This paper presents a risk–budgeted monitor with a control framework that certifies safety for autonomous driving. In this process, a sliding window is proposed to monitor for insufficient barrier residuals or nonzero tail risk, ensuring system safety. When the safety margin deteriorates, it triggers switching the safety constraint from a performance-based relaxed-control barrier function (R-CBF) to a conservative conditional value at risk (CVaR-CBF) to address the safety concern. This switching is governed by two real–time triggers: Feasibility–Triggered (FT) and Quality–Triggered (QT) conditions. In the FT condition, if the R-CBF constraint becomes infeasible or yields a suboptimal solution, the risk monitor triggers the use of the CVaR constraints for the controller. In the QT condition, the risk monitor observes the safety margin of the R-CBF solution at every step, regardless of feasibility. If it falls below the safety margin, the safety filter switches to the CVaR-CBF constraints.
The proposed framework is evaluated using a model predictive controller (MPC) for autonomous driving in the presence of autonomous vehicle (AV) localization noise and obstacle position uncertainties. Multiple AV-pedestrian interaction scenarios are considered, with 1,500 Monte Carlo runs conducted for all scenarios. In the most challenging setting with pedestrian detection uncertainty of 5m, the proposed framework achieves a 94-96% success rate of not colliding with the pedestrians over 300 trials while maintaining the lowest mean cross–track error (CTE=3.2-3.6 m) to the reference path. The reduced CTE indicates faster trajectory recovery after obstacle avoidance, demonstrating a balance between safety and performance.
I Introduction
Safety, despite uncertainties, is a crucial aspect when developing controllers for Automated vehicles (AVs). These uncertainties drastically increase the computational demand, and the problem becomes challenging when the controller feedback frequency typically needs to operate between 20205050 Hz [1].
Control Barrier Functions (CBFs) in quadratic programming (QP) formulation is one of the most fundamental and widely used strategies in the real-time implementation of obstacle-avoidance, safety-guaranteeing controllers. However, through the principle of forward invariance, CBFs can guarantee safety only when the QP is feasible and optimal [2, 3, 4]. To improve deployability, prior work has explored relaxed CBFs (R-CBF) with slack penalties [5], the interplay between input limits and barrier constraints [6, 7], higher-order CBFs to solve the high-degree dynamic system [8], and learning-driven tuning of controller and safety parameters [9]. Under environmental uncertainties, stochastic CBFs [10, 11, 12] and CVaR-based CBFs [13, 14] address tail events (like unpredictable pedestrians) at additional computational cost and conservative behaviors.
Figure 1: Proposed framework with MPC nominal control and a risk–budget monitor for safety–driven controller switching.
Several works acknowledge that input and safety constraints can render the inner QP infeasible and pursue mitigation via parameter tuning or adaptation [9, 14]. Others address feasibility through hierarchical optimization or learning-based augmentation of feasible sets [15]. While these approaches improve feasibility, they often trade off computational budget or require substantial data collection to bias the controller toward feasible solutions, and they may not directly quantify safety when infeasibility persists under real-time limits.
This paper begins from a practical perspective, suggesting that infeasibility does not necessarily imply a loss of safety. Infeasible or suboptimal steps may arise from calculation time limitations or input constraints rather than imminent safety violations. Focusing on the safety margin with respect to the system, we propose a risk-budgeted monitor that evaluates the safety quality of the control action over a desired window length. Instead of penalizing a single suboptimal or infeasible control action, we now look over a window to certify safe behavior of the AV. This allows the system to recover to a safe state despite occasional undesired control actions that might have occurred due to computational complexity or input limitations. Inferring that monitoring the safety margin does not make the system safer, we utilize the risk monitor warning as a switch signal to determine when to use a conservative safety filter. In short, we run a relaxed CBF-QP by default for trajectory tracking. When the risk monitor detects a decline in safety quality, the filter will switch to a conservative CVaR–CBF to bound the tail risk under uncertainty. The overall architecture is shown in Fig. 1. This work makes the following key contributions.
  1. 1.
    We formalize a risk–budget monitor that provides windowed safety guarantees over a finite horizon.
  2. 2.
    We develop a monitor–driven switching mechanism between a performance–oriented CBF–QP and a conservative CVaR–CBF, enabling real-time navigation with bounded tail risk.
  3. 3.
    We theoretically and experimentally show that the risk budgeting permits controlled, transient infeasibility while preserving certified finite-horizon safety.
II Related Work
Recent work tackles the interplay between infeasibility and safety constraints from complementary angles. Ma et al. embed an exponential CBF QP as a differentiable layer so that the safety shaping is learned rather than hand-tuned [16]. A penalized slack is used only during training to keep the QP solvable while back-propagating. Manda et al. instead learn the barrier itself so that the control-invariant safe set is enlarged under actuation limits, lowering the likelihood that the safety-filter QP becomes empty online [17]; the trade-off is scalability and expressivity at the expense of limited formal robustness guarantees. Overall, these data-driven approaches address infeasibility by learning either the shaping, the tuning, or the certificate, while keeping safety as a hard constraint. However, their effectiveness can be scenario dependent and operation design domain limitation, requiring representative data and careful validation before broad deployment.
Ahmadi et al. introduce CVaR barrier functions (CVaR-BFs) for discrete-time systems that enforce safety on the distributional tail [13]. For dynamic crowded environments, Wang et al. make the feasibility–safety explicit that smaller risk levels contract the feasible set, whereas larger risk levels expand it at the cost of conservatism; they adapt the CVaR confidence level online to the minimum required for feasibility and augment with a dynamic-zone CBF that enlarges the virtual safety radius under risky relative motion, preserving the intended probabilistic guarantee [14]. Across these works, infeasibility is mitigated by learning risk parameters and exploiting local feasibility of the stepwise optimization, but universal feasibility certificates under tight actuation and aggressive interactions remain open, with computational load. [15]. In both cases, safety remains the top priority—either by fencing off infeasible states or by guaranteeing a feasible least-violation solution when perfect invariance is impossible.
The chapter of the paper is organized as follows. We first review standard and relaxed CBF formulations, then define stochastic safe sets and introduce the relaxed CVaR–CBF formulation for real-time implementation. Section IV formalizes the sliding-window risk-budgeted monitor for safety risk, and Section V details the triggering logic for mode switching. We conclude with numerical experiments and a discussion of results.
III Preliminaries
Consider a nonlinear continuous time, control-affine system given by:
$\dot{\bm{x}}=f(\bm{x})+g(\bm{x})\bm{u},$ (1)
where $\bm{x}\in\mathcal{D}\subset\mathbb{R}^{n}$ and $\bm{u}\in\mathcal{U}\subset\mathbb{R}^{m}$ represent the system’s state and input. $\mathcal{U}$ is the set of admissible inputs. The functions $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n},\quad$ and $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times m}$ are locally Lipschitz continuous.
The safe set $\mathcal{S}=\{\bm{x}:h(\bm{x})\geq 0\}$ is the zero-superlevel set of a function $h:D\subset\mathbb{R}^{n}\rightarrow\mathbb{R}$, and the system is guaranteed to be safe if the control input satisfies:
$L_{f}h(\bm{x})+L_{g}h(\bm{x})\bm{u}+\kappa h(\bm{x})\ \geq\ 0,$ (2)
where $\kappa$ is a user-defined positive constant that determines the minimum decay rate of $h(\bm{x})$[18]. The dynamics of the barrier function is given by:
$\dot{h}(\bm{x},\bm{u})=L_{f}h(\bm{x})+L_{g}h(\bm{x})\bm{u}.$ (3)
Let a safety residual be defined as $r(\bm{x},\bm{u})$:
$r(\bm{x},\bm{u}):=L_{f}h(\bm{x})+L_{g}h(\bm{x})\bm{u}+\kappa h(\bm{x}).$ (4)
For implementing the controller in a real-time system like autonomous vehicle (AV), the continuous-time system is discretized using a zero-order hold (ZOH)
$\bm{u}(t)=u_{k},\quad\forall t\in[t_{k},t_{k+1}),$ (5)
where $u_{k}$ is the constant control input between the sampling time $t_{k}$ and $t_{k+1}$. The sampling has a uniform sampling time interval $T_{s}$ [19]. Thus, as in (5), the barrier dynamics in (3) satisfy
$\dot{h}(\bm{x},u_{k})=-\kappa h(\bm{x},u_{k})+r(\bm{x},u_{k}).$ (6)
The sample-wise lower bound is given by
$r_{k}=\inf_{t\in[t_{k},t_{k+1})}r\big(\bm{x},u_{k}\big)$ (7)
$\qquad(\text{for small }T_{s},\ r_{k}\approx r(\bm{x},u_{k})).$
Since $\bm{x}$ is a time varying function, we can rewrite (6) as:
$\dot{h}(t)\geq-\kappa h(t)+r_{k}.$ (8)
Solving this linear differential inequality yields the one-step comparison given by:
$h_{k+1}\ \geq\ \mu\,h_{k}+c\,r_{k},$ (9)
where $\mu=e^{-\kappa T_{s}}\in(0,1)$ and $c=\frac{1-\mu}{\kappa}$. This is a sufficient constraint on the next-step state $h_{k+1}$ given $(h_{k},u_{k})$, ie, it is a guaranteed lower bound rather than an exact point prediction.
Remark 1
In our experiments, we choose $T_{s}$ to be sufficiently small so that the one-step bound (9) remains valid, while minimizing the effect of sampled-data mismatch.
III-A Relaxed CBF (R-CBF)
The theory thus far yields a safe control input $u_{k}$ while a nominal controller input $u_{k}^{\text{nom}}$ is provided by the MPC in our implementation at time $k$. Thus, an optimal safe controller is implemented in a QP formulation using a CBF-based safety filter with a nonnegative slack. The slack addresses infeasibility due to actuator limitations as shown in  [18]:
$\min_{u_{k}\in\mathcal{U},\;\nu_{k}\geq 0}$ $\tfrac{1}{2}\|u_{k}-u_{k}^{\text{nom}}\|^{2}\;+\;\rho_{\nu}\,\nu_{k}^{2}$ (10)
s.t. $L_{f}h(x_{k})+L_{g}h(x_{k})\,u_{k}+\kappa\,h(x_{k})\;\geq\;-\,\nu_{k}.$
The slack $\nu_{k}>0$ quantifies the instantaneous relaxation of the barrier constraint and is penalized by $\rho_{\nu}>0$. A smaller penalty admits larger (more aggressive and performance-seeking) relaxations, while a larger penalty on $\nu_{k}$ yields more conservative (safety-biased) behavior. In Section IV, we demonstrate how to utilize $\nu_{k}$ and the residual margins $r_{k}$ as supervisory signals that guide a balanced control framework between performance and safety.
III-B Stochastic Safe Sets
Given the inherent uncertainties in dynamic obstacle positions and vehicle localization, a safe set for the AV cannot be computed deterministically. To address this challenge, we transition to a stochastic formulation where safety is evaluated in terms of probabilistic risk. The obstacle state is sampled $P$ times from its distribution, denoted as $\bm{x}^{o,p}\in\mathcal{O}\subset\mathbb{R}^{n}$, where $p=1,\dots,P$. Similarly, the vehicle state is sampled $Q$ times, denoted as $\bm{x}^{q}\in\mathcal{D}\subset\mathbb{R}^{n}$ where $q=1,\dots,Q$. These independent samples are then combined to form a total of $S=P\times Q$ stochastic safe set. To represent each unique vehicle-obstacle pair, we use a double index $(p,q)$. The CBF for each safe set at time k is defined as:
$h_{k}^{i}$ $=\left\|{\mathbf{x}}_{k}^{v,q}-{\mathbf{x}}_{k}^{o,p}\right\|_{2}-D_{s},$ (11a)
where $i=1,\dots,S$ is the total number of pair $(p,q)$ and $D_{s}$ represents the minimum safe distance. The safety constraints $h_{k+1}$ and the residual $r_{k}$ can be written in stochastic formulation as:
$h_{k+1}^{i}$ $\geq\mu h_{k}^{i}+cr_{k}^{i},$ (11b)
$r_{k}^{i}(u_{k})$ $=L_{f}h_{k}^{i}+L_{g}h_{k}^{i}u_{k}+\kappa h_{k}^{i}.$ (11c)
To align the goal of maximizing safety with a standard minimization problem, we define a loss function $Z_{k}^{i}(u_{k})$ as the negative of the safety residual:
$Z_{k}^{i}(u_{k}):=-r_{k}^{i}(u_{k}).$ (12)
This definition creates an intuitive relationship: a highly safe outcome (large positive $r_{k}^{i}$) corresponds to a very low loss (negative $Z_{k}^{i}$), while an unsafe outcome (negative $r_{k}^{i}$) results in a high, positive loss. This transformation turns the safety problem into a convex, sample-average tail control problem.
III-C Risk-Aware Safety Critical Control Formulation
III-C1 Probability Constraints
Due to the stochasticity in obstacle and vehicle positions, the optimization of the loss function is considered a chance-constrained optimization problem that accommodates uncertainties, as given by.
$\mathbb{P}\left(Z_{k+1}^{i}\leq 0\right)\geq\epsilon,$ (13)
where $\epsilon\in(0,1)$ represents the desired confidence level. This constraint ensures that the probability of the system remaining in a safe state (defined by a non-positive loss, $Z_{k+1}^{i}\leq 0$) is at least $\epsilon$. From (13), we can formulate the probability constraints to a standard measure of risk using Value-at-Risk (VaR). The VaR of the loss function at the confidence level $\epsilon$ can be defined as:
$\text{VaR}_{\epsilon}(Z_{k+1}^{i}):=\inf_{\gamma\in\mathbb{R}}\left\{\gamma\;\middle|\;\mathbb{P}(Z_{k+1}^{i}\leq\gamma)\geq\epsilon\right\},$ (14)
where $\gamma$ represents the candidate loss threshold being evaluated. The relationship between this candidate threshold $\gamma$ and the safety boundary is the key to this formulation. The initial probabilistic constraint in (13) establishes a specific requirement that the system must satisfy. Comparing the safety boundary to the VaR definition, $\mathbb{P}(Z_{k+1}^{i}\leq\gamma)\geq\epsilon$, it is evident that $\gamma=0$ must be one of the values that satisfy the safety condition.
This leads to the final, tractable constraint:
$\text{VaR}_{\epsilon}(Z_{k+1}^{i})\leq 0.$ (15)
This means that satisfying the constraint $\text{VaR}_{\epsilon}(Z_{k+1}^{i})\leq 0$ provides a probabilistic guarantee of safety. It ensures the system will not enter an unsafe state under the uncertainties considered with a confidence level of $\epsilon$.
However, VaR provides no information about the magnitude of the expected loss within the $(1-\epsilon)$ risk tail of the distribution. Once an outcome falls into this tail, the loss value could be close to zero or a very large number, which determines the severity of the safety violation. In turn, Conditional Value-at-Risk (CVaR) is introduced to address this limitation.
Definition 1 (Conditional Value-at-Risk). The CVaR at a confidence level $\epsilon$ is the expected loss in the $(1-\epsilon)$ tail of the distribution ($Z_{k+1}^{i}$), which occurs when the loss exceeds threshold $\text{VaR}_{\epsilon}$ [20].
$\text{CVaR}\epsilon(Z_{k+1}^{i}):=\mathbb{E}\left[Z_{k+1}^{i}\middle|Z_{k+1}^{i}\geq\text{VaR}_{\epsilon}(Z_{k+1}^{i})\right].$ (16)
From [21], the (16) can be rewritten as:
$\text{CVaR}_{\epsilon}(Z_{k+1}^{i})=\inf_{\gamma\in\mathbb{R}}\left\{\gamma+\frac{1}{1-\epsilon}\,\mathbb{E}\!\left[(Z_{k+1}^{i}-\gamma)_{+}\right]\right\},$ (17)
where $(x)_{+}:=\max(0,x)$ is the positive part function. The infimum is achieved when $\gamma$ is equal to $\text{VaR}_{\epsilon}(Z_{k+1}^{i})$. The sample-based CVaR of the loss $Z_{k}^{i}$ can be found by solving for the optimal risk threshold $\gamma$.
In practice, finding a control solution that completely eliminates risk is not always feasible. Therefore, we introduce a slack variable $\nu_{k}$ to represent the accepted level of risk, and refer to the resulting method as the Relaxed CVaR–CBF. The safety filter then minimally modifies a nominal command $u_{k}^{\text{nom}}$ by solving the following quadratic programming (QP) problem:
[…]
$\text{s.t.}\quad\text{CVaR}\epsilon(Z_{k+1}^{i})\leq\nu_{k},$ $0\leq\nu_{k}\leq\overline{\nu},$
$r_{k}^{i}(u_{k})\ \geq\ -\,\nu,\quad i=1,\dots,S,$
where $\nu_{k}$ represents the upper bound of the tail severity, while $\overline{\nu}$ acts as a risk cap, limiting the maximum allowable risk for any single step. The window-level safety certificate determines the risk cap’s maximum allowable magnitude to guarantee safety over a finite time horizon, as discussed in IV-B.The coefficient $\rho_{\nu}>0$ is a penalty weight, and the term $\rho_{\nu}\nu_{k}^{2}$ in the cost function penalizes any non-zero risk. By allowing $\nu_{k}$ to be non-zero, this formulation makes the controller more robust by preventing infeasibility in challenging situations, while $\nu_{k}$ provides a direct, quantifiable measure of the risk taken in that step.
IV Risk-Budget Monitor
As the optimization in (18) allows for non-zero risk, a balance can be struck between performance, safety, and the prevention of solution infeasibility. Furthermore, a single infeasible step or a momentary non-zero risk might not immediately lead the AV into an unsafe area, especially when the system operates with a safety buffer. To manage this trade-off effectively, this section defines the criteria for bad step detection and the mechanism of the risk window monitor.
IV-A Bad Step detection
Before defining the bad step condition, we must first define the parameter $\delta$, which represents a pre-defined safety margin required for a good step. and it should be a positive number $\delta>0$. A step is considered good only if the residual $r_{k}$ is greater than or equal to $\delta$, effectively depositing a guaranteed safety allowance into the system.
Accordingly, a step is classified as a bad step if it meets either of the following two conditions:
  • Condition I: $\nu_{k}>\bar{\nu}$
    The variable $\nu_{k}$ quantifies the controller’s acceptance of finding a feasible solution, and $\bar{\nu}$ is the worst allowance risk. If $\nu_{k}$ is greater than $\bar{\nu}$, it signifies that the chosen control action has exceeded the maximum allowable risk. Therefore, the step is classified as a bad step.
  • Condition II: $r_{k}<\delta$
    This condition flags a step as bad if the residual in R-CBF fails to meet this minimum safety margin $\delta$ required for a good step. This acts as a conservative check; a step is considered bad even if the overall statistical risk is zero ($\nu_{k}=0$), ensuring that any single foreseeable negative outcome is accounted for.
The bad step indicator $b_{k}$ at time step $k$ is defined from the conditions as:
$b_{k}=\mathbf{1}\{\nu_{k}>\bar{\nu}\}\ \lor\ \mathbf{1}\{r_{\min,k}<\delta\},\quad b_{k}\in\{0,1\},$ (19)
i.e., $b_{k}=1$ denotes a bad step, otherwise the step is good.
IV-B Sliding Window Risk-Budgeted Monitor
In the previous section, we defined the conditions for classifying a control action as a bad step in (19). At each time step, the system applies these criteria, effectively abstracting the complexity of the many individual samples into a deterministic good or bad classification.
This classification is tracked by a risk monitor that operates over a sliding window of length $W$, ensuring that no more than $M$ bad steps, which withdraw from the system’s safety allowance within the window. However, the allowable ratio of bad steps ($M/W$) is not arbitrary; it is fundamentally linked to the good-step safety margin, $\delta$, and the maximum permissible risk, $\overline{\nu}$. This section derives the precise mathematical relationship that governs these parameters, proving the safety of the monitor.
The analysis begins with the discrete single-step inequality from (9), which is recursively applied over a window of length $W$. This process evaluates the cumulative effect of a sequence of control actions, governed by the rule that no more than $M$ bad steps can occur.
By unrolling the inequality for $W$ steps, we obtain the following expression for the lower bound of $h_{k}$:
$h_{k}\geq\mu^{W}h_{k}+c\sum_{j=0}^{W-1}\mu^{W-1-j}v_{k-j}.$ (20)
Here, $v_{k-j}$ is a variable representing the worst-case lower bound of the residual $r_{k-j}$ used in the proof. Thus, $v_{k-j}$ takes a value from the set $\{\delta,-\overline{\nu}\}$. This is because a good step is guaranteed to have a residual of at least $\delta$ (in section IV-A), while a bad step has its residual lower-bounded by $\overline{\nu}$. To guarantee safety ($h_{k}\geq 0$), we must find a lower bound for this expression. The core challenge lies in finding the worst-case arrangement of the $M$ bad steps in the summation. The next lemma establishes that the worst case occurs at the end of the step. Consequently, we can apply the rearrangement lemma to determine the parameters in this worst-case configuration.
Lemma 1 (Rearrangement Lemma)
Let the weights $w_{j}:=\mu^{W-1-j}$ for $j=0,\dots,W-1$. Since $\mu\in(0,1)$, these weights are monotonically increasing: $0<w_{0}<w_{1}<\dots<w_{W-1}$. Let $V$ be the set of sequences $\{v_{j}\}_{j=0}^{W-1}$ with the worst case $M$ entries equal to $\overline{\nu}$ and WMW-M entries equal to $\delta$. The weighted sum $S(v):=\sum_{j=0}^{W-1}w_{j}v_{j}$ is minimized over $v\in V$ by assigning the $M$ negative values $(-\overline{\nu})$ to the largest weights and the WMW-M positive values $(+\delta)$ to the smallest weights.
Proof:
Consider any pair of indices $i<j$, which implies $w_{i}\leq w_{j}$. Let’s assume an arrangement where $v_{i}=\delta$ and $v_{j}=-\overline{\nu}$. The contribution from this pair to the sum is $w_{i}\delta+w_{j}(-\overline{\nu})$. By swapping the values: $v^{\prime}_{i}=-\overline{\nu}$ and $v^{\prime}_{j}=\delta$. The new contribution is $w_{i}(-\overline{\nu})+w_{j}\delta$. The difference between the original sum and swapped sum is:
$(w_{i}\delta-w_{j}\overline{\nu})-(-w_{i}\overline{\nu}+w_{j}\delta)$ $=w_{i}(\delta+\overline{\nu})-w_{j}(\overline{\nu}+\delta)$
$=(w_{i}-w_{j})(\delta+\overline{\nu}).$
Since $w_{i}\leq w_{j}$ and $\delta,\overline{\nu}\geq 0$, the difference $(w_{i}-w_{j})(\delta+\overline{\nu})$ is less than or equal to 0. This means the sum is minimized when the smaller value ($\overline{\nu}$) is paired with the larger weight ($w_{j}$). By repeatedly applying such pairwise swaps, we can prove that the minimum sum is achieved when all negative values are assigned to the largest weights. ∎
Applying the Rearrangement Lemma to the summation in Equation (20), we can establish the definitive worst-case lower bound for $h_{k}$. The minimum is achieved when the $M$ bad steps are assigned to the terms with the largest weights (corresponding to indices $j=W-M,\dots,W-1$) and the WMW-M good steps are assigned to the terms with the smallest weights (indices $j=0,\dots,W-M-1$).
This splits the single summation into two parts. For notational clarity, we then re-index the sums. By substituting the summation index via the transformation l=W1jl=W-1-j, the weights $\mu^{W-1-j}$ simply become $\mu^{l}$. This transformation maps the range for good steps ($j=0,\dots,W-M-1$) to $l=M,\dots,W-1$, and the range for bad steps ($j=W-M,\dots,W-1$) to $l=0,\dots,M-1$. This substitution yields the final lower bound:
$h_{k}\geq\mu^{W}h_{k}+c\left[\delta\sum_{l=M}^{W-1}\mu^{l}-\overline{\nu}\sum_{l=0}^{M-1}\mu^{l}\right].$ (21)
This inequality represents the guaranteed system state under the worst possible sequence of actions and serves as the foundation for our main theorem.
Theorem 1 (Window-Level Safety Certificate)
Assume the one-step comparison inequality (9) holds with $0<\mu<1$, $c>0$. In any sliding window of length $W$, there are at most $M$ bad steps (with $r\geq-\overline{\nu}$) and WMW\!-\!M good steps (with $r\geq\delta$). Moreover, each window starts from a safe boundary state $h_{k}\geq 0$. Then the terminal state is safe ($h_{k}\geq 0$) whenever
$\mu^{M}(1-\mu^{W-M})\,\delta\ \geq\ (1-\mu^{M})\,\overline{\nu},$ (22)
where $\mu=e^{-\kappa T_{s}}$.
Proof:
To ensure safety, we require that the lower bound of $h_{k}$ remains non-negative. Starting from the worst-case inequality established in (21) and conservatively dropping the $\mu^{W}h_{k}\geq 0$ term, the condition becomes:
$c\left[\delta\sum_{l=M}^{W-1}\mu^{l}-\overline{\nu}\sum_{l=0}^{M-1}\mu^{l}\right]\geq 0.$ (23)
Evaluating the geometric sums gives:
$\delta\frac{\mu^{M}(1-\mu^{W-M})}{1-\mu}\geq\overline{\nu}\frac{1-\mu^{M}}{1-\mu}.$ (24)
Since $c>0$ and $0<\mu<1$ so that $(1-\mu)>0$, multiplying both sides of (23) by $(1-\mu)/c$ preserves the inequality direction and gives the certificate in (22). ∎
Rather than forbidding any instantaneous violation, the certificate caps how many bad steps and how severe they can be via $(W,M,\delta,\overline{\nu})$, thereby guaranteeing $h_{k}\geq 0$ at every window boundary. Once the $W,M,\delta$ has been designed based on the system requirement, we can get $\overline{\nu}$ from the equation below:
$\overline{\nu}\leq\frac{\mu^{M}(1-\mu^{W-M})}{1-\mu^{M}}\delta.$ (25)
V Risk-Switch for Probabilistic Safety
Building on the risk monitor in Section IV, we formalize a sliding-window supervisor that certifies safety at the window scale and coordinates switching from the performance-seeking R-CBF filter to a conservative RC-CBF safety filter when risk accumulates. The cumulative number of bad steps $m_{k}$ in the most recent $W$-step window is updated by
$m_{k}\;=\;m_{k-1}\;+\;b_{k}\;-\;b_{k-W},$ (26)
with $m_{k}\in\{0,1,\dots,W\}$.
V-A Feasibility-Triggered CVaR-CBF (FT-C-CBF)
Let $\mathcal{A}$ denote the feasible set of the R-CBF filter at time $k$, and let $a_{k}=\mathbf{1}\{(u_{k},\nu_{k})\in\mathcal{F}\}$ be its feasibility flag. Under FT, feasibility is the primary guard, while the sliding-window counter enforces a budget on accumulated risk:
If the R-CBF problem is infeasible at time $k$ ($a_{k}=0$), the monitor updates $b_{k}$ and $m_{k}$ via (26). When the cumulative count reaches the budget $M$ (i.e., $m_{k}\geq M$), the monitor switches from the R-CBF filter to the CVaR-CBF controller.
$\eta_{k}=\begin{cases}1,&\text{if }(a_{k}=0)\ \land\ (m_{k}\geq M),\\ 0,&\text{if }(a_{k}=1)\ \land\ (m_{k}<M),\end{cases}$
with $\eta_{k}\in\{0,1\}$ the controller mode (0: R-CBF, $1$: CVaR-CBF). The controller $\eta_{k}$ will be implemented at time $k$.
V-B Quality-Triggered CVaR-CBF (QT-C-CBF)
In the QT mechanism, every step is vetted after solving the R-CBF filter, irrespective of feasibility, the risk monitor computes $b_{k}$ and updates the sliding-window count $m_{k}$. Switching depends solely on whether the window-level risk budget is exceeded.
$\eta_{k}=\begin{cases}1,&\text{if }m_{k}\geq M,\\ 0,&\text{if }m_{k}<M,\end{cases}$
Thus, even when the R-CBF problem is feasible, a step flagged as bad by the monitor contributes to $m_{k}$; once the cumulative count reaches the budget $M$, the supervisor switches from the R-CBF filter to the CVaR controller.
Figure 2: Success rate versus mean cross–track error (CTE) relative to the reference trajectory across 1,500 runs.
VI Simulations
VI-A Experimental Setup
VI-A1 Environment
We evaluate a kinematic vehicle under an MPC controller, following the experimental setup of [1]. The reference trajectory is converted from a GPS path collected from the experimental autonomous vehicle reported in [1] and validated.
VI-A2 Safety Framework Configuration
The minimum safety distance set to $D_{s}$=3m, and $\kappa=1$, and the sliding window risk monitor is configured with a window size of W=5, a maximum of M=1 bad step, and a good step safety margin $\delta=1$m. Based on these parameters, the risk cap is calculated using (25) to be $\overline{\nu}$= 3.8.
The Conditional Value-at-Risk (CVaR) uses a confidence level of $\epsilon$=95% and a loss threshold of $\gamma$=0, consistent with the definition in (17). To simulate realistic localization uncertainty, the vehicle’s position is modeled as a Gaussian distribution. At each timestep, Q=10 samples are drawn from the distribution, which is centered on the vehicle’s measured position and has a standard deviation $\sigma_{v}=0.1$, to represent the set of possible true locations.
VI-A3 Obstacle Uncertainty Modeling
To account for obstacle movement and perception errors, we model uncertainty in the obstacle’s state. At each timestep $k$, upon receiving an obstacle’s measured position $\bm{x}_{m,k}^{o}$, we generate P=10 samples to represent its possible locations $\bm{x}_{k}^{o,p}$. These samples are uniformly distributed within a square bounding box of half width $\sigma_{o}\in\{1,2,3,4,5\}$m centered on the measured position. This setup simulates a perception detection that provides a bounded box estimate of the obstacle’s location, with the chosen parameters representing an uncertainty scenario.
$\bm{x}_{k}^{o,p}=\bm{x}_{m,k}^{o}+\Delta\bm{x}_{k}^{o,p},$ (27)
$\Delta\bm{x}_{k}^{o,p}\sim\mathcal{U}\left(\left[-{\sigma_{o}},{\sigma_{o}}\right]^{2}\right).$
VI-A4 Performance Metrics
To evaluate the controller’s performance, we define five metrics to aggregate performance over all test cases. Some of our aggregate metrics are adapted from the methodology presented in [14].
To assess overall performance across all $N_{t}$ test cases, we calculate the following metrics. Note that for metrics other than success rates, the averages are computed only over the set of successful runs, $N_{s}$.
  • Success Rate (SR): A run is successful if the minimum vehicle pedestrian distance along the trajectory always exceeds the required clearance $d_{\min}=2.8\,\mathrm{m}$ (vehicle radius $1.8\,\mathrm{m}$ + pedestrian radius $1.0\,\mathrm{m}$). The success rate is the fraction of successful runs: $\mathrm{SR}=N_{s}/N_{t}$.
  • Minimum Distance to Pedestrian (MDP): For each successful run, let $d_{m,i}$ be the minimum distance to the nearest pedestrian while avoiding; we report the mean over successful runs, $MDP=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}d_{m,i}$. An ideal clearance of at least $5.8\,\mathrm{m}$ is required (vehicle radius $1.8\,\mathrm{m}$ + pedestrian radius $1.0\,\mathrm{m}$ + safety buffer $3.0\,\mathrm{m}$). Values $\geq 5.8\,\mathrm{m}$ satisfy the safety margin; smaller values indicate the controller is performance-based to track the reference path as close as possible.
  • Infeasibility Rate (IR): The proportion of optimization steps that failed to yield a feasible solution, calculated as $IR_{i}=n_{i}/n_{t}$, where $n_{i}$ is the number of infeasible steps and $n_{t}$ is the total number of optimization steps. The mean infeasibility solution rate across all successful runs, $IR=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}IR_{i}$.
  • Calculation Time (CT): The mean calculation time for each step in all successful runs, $CT=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}T_{i}$.
  • Cross-Track Error (CTE): To evaluate path-following performance, we calculate the deviation from the desired path. We record the Average CTE ($E_{i}$), calculated as the mean of all CTE measurements $e_{k}$ during the run. The mean of the average CTE values from all successful runs, $CTE=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}E_{i}$.
VI-B Comparative Analysis
In the comparative evaluation, MPC serves as the nominal controller and is composed of the following safety filters:
  1. 1.
    Relaxed CBF (R-CBF): We enforce the CBF constraint via the relaxed formulation in (10) over the entire trajectory, using the raw, noisy pedestrian measurements as inputs to the filter (i.e., no stochastic pedestrian and vehicle states in the controller).
  2. 2.
    CVaR–CBF (hard tail constraint) (C-CBF): The CVaR surrogate in (17) and remove risk slack by fixing $\nu_{k}\equiv 0$ in (18). This yields a strict tail-risk constraint, $E^{+}(u_{k})\leq 0$, with no per-step relaxation.
  3. 3.
[…]
  4. 4.
    Relaxed CVaR–CBF (RC-CBF): We use the relaxed CVaR–CBF defined by (17) and (18) throughout the trajectory, allowing $\nu_{k}\geq 0$ and no risk monitor.
The results are evaluated with 1,500 Monte–Carlo runs under vehicle localization noise $\sigma_{v}=0.1$ and pedestrian position noise $\sigma_{o}\in\{1,2,3,4,5\}\,\mathrm{m}$. We consider three different pedestrian interaction scenarios: (i) one pedestrian performing a jaywalk across the AV’s path, (ii) two pedestrians, and (iii) three pedestrians, with 1,500 total runs. Figure 2 organizes results by pedestrian number. The figure reports the detection bounding-box uncertainty; the left y-axis shows success rate (%), and the right y-axis shows the mean cross–track error (CTE) computed over successful runs. Across uncertainty levels, the proposed framework FT and QT triggers sustain high success rates while maintaining comparatively low CTE relative to the baselines.
Figure 3: (a) AV trajectory comparison for R-CBF, RC-CBF, FT-C-CBF, and QT-C-CBF under $\sigma_{v}=0.1,\sigma_{o}=5$ with three pedestrians. (b) Stochastic safe set: sampling-based barrier values $h(x)$ with sample mean (solid) and sampled range (shaded); shading appears when CVaR control is active. (c) Deterministic safe set: $h(x)$ evaluated at true states of vehicle and pedestrians without sampling, highlighting clearance relative to $h(x)=0$.
Most Challenging Scenario
The table summarizes all pedestrian interaction scenarios with $\sigma_{v}=0.1$ and $\sigma_{o}=5\,\mathrm{m}$. Table I indicates that the conservative baselines C-CBF and RC-CBF achieve the highest success rates (SR $98\%,97\%$), but maintain very large minimum distances to pedestrians (MDP $11.3,12\,\mathrm{m}$), almost twice the required $5.8\,\mathrm{m}$. This conservatism is accompanied by larger cross-track error (CTE $4.2,4.3\,\mathrm{m}$) and notable computation time (CT $56.2,75.3\,\mathrm{ms}$). In contrast, the performance-based R-CBF has the lowest SR ($67\%$) with the smallest MDP ($6.3\,\mathrm{m}$) and the smallest CTE ($3.9\,\mathrm{m}$), reflecting aggressive behaviors that occasionally fails. Our design goal is to preserve C-CBF level safety while driving MDP and CTE toward the R-CBF baseline.
[…]
Figure 3 shows one of the test runs results, in (a) four control trajectories has been provided: R-CBF collides between the second and third pedestrians; RC-CBF yields the farthest detours; the proposed FT-C-CBF and QT-C-CBF maintain moderate clearance and recover to the reference faster than RC-CBF. Panel (b) plots $h(x)$ with sample mean (solid) and sampled range (shaded). Two effects are visible: (i) FT limits the duration of low–$h$ excursions by switching when feasibility is threatened, while QT prevents deep dips by switching earlier on quality degradation. Panel (c) plots the barrier value $h(x)$ along the true state of the vehicle and pedestrian trajectory, reflecting the true safe boundary situation. The R-CBF drops below $h=0$, indicating a collision event. In contrast, C-CBF (and RC-CBF) remain above the boundary safe but conservative. The FT and QT triggers skim the boundary and rapidly recover, consistent with their moderate minimum distance (MDP) and low CTE achieved by switching between R-CBF and CVaR-CBF.
TABLE I: The comparison is under $\sigma_{v}=0.1,\sigma_{o}=5$ with three pedestrian interactions, with total 300 runs.
Method SR (%)
MDP (m)
($\geq$ 5.8 m)
IR (%) CT (ms) CTE (m)
R-CBF 67% 6.3 11.9 % 22.4 3.9
C-CBF 98% 11.3 15.2% 56.2 4.2
AC-CBF 84% 7.33 17.8% 76.9 5.0
RC-CBF 97% 12 15.3% 75.3 4.3
FT-C-CBF 94% 7.1 15.8% 24.2 3.2
QT-C-CBF 96% 8.3 16.1% 25.7 3.6
TABLE II: 300 runs across scenarios; $\delta\in\{0.1,1,2\}$.
$\delta=0.1$ $\delta=1$ $\delta=2$
FT QT FT QT FT QT
SR 94% 98% 94% 97% 93% 96%
MDP (m) 7.26 9.71 7.05 9.04 7.2 8.8
CTE (m) 3.5 4.14 3.5 4.1 3.5 3.9
CVaR rate 0.92% 2.61% 0.96% 2.70% 1.04% 2.93%
VI-C Delta Comparison
With the risk window fixed at $(W,M)=(5,1)$, we sweep the safety margin parameter $\delta\in\{0.1,1,2\}$ across the three pedestrian scenarios, the risk cap is calculated using (25) to be $\overline{\nu}$= 0.38, 3.8, 7.6 corresponding to the $\delta$ in (Table II). Because $\delta$ is coupled to the admissible risk slack $\bar{\nu}$, varying $\delta$ primarily shifts the threshold that triggers the conservative CVaR controller, without otherwise altering the nominal MPC objective. In addition to SR, MDP, and CTE, we report the CVaR rate, defined as the percentage of steps at which the CVaR controller is engaged. Consistent with design, the activation rate increases monotonically with $\delta$ for both FT and QT switching, reflecting a larger safety margin that induces earlier and more frequent activations.
FT vs. QT
From Table II, QT has a consistently higher SR than FT at each tested $\delta$. This follows from the QT design: the risk–budget monitor is enforced at every step, so trajectories that are optimally feasible but do not meet the desired safety margin. QT therefore triggers the $\mathrm{CVaR}$–CBF more often than FT, which is reflected in the higher $\mathrm{CVaR}$ activation rates reported in Table II.
VII Conclusion
We presented a control framework with a risk monitor to switch the safety filter between the performance CBF and conservative CVAR in autonomous driving scenarios. Monte–Carlo evaluations with localization and obstacle state uncertainty show that using the QT option in the control framework has a high success rate ($\geq\!95\%$) and maintains low CTE showing the trajectory following performance while avoiding the pedestrians. When safety quality is high, the risk monitor allow the controller in its performance mode; as the margin degrades, the monitor reliably triggers the CVaR controller to preserve safety throughout the trajectory. The proposed framework achieves a performance safety balanced with risk-budgeted guarantees and potential computation compatible with on–board control loops. As next steps, we will deploy the framework under the proposed risk-budgeted monitor on our autonomous vehicle platform for hardware–in–the–loop and on–road validation.